# Strategic Informed Trading and the Value of Private Information

### v1

@@ Appendix E Proofs from Section 3
Next define the function (c.f. (34))
$Q(p,y)=\frac{\kappa p(1+p)+y^{2}}{\kappa(1+p)(1+2y)}$
so that
@@
At this point, if one plugs these values into the right side of (93) and takes a common denominator, the numerator is a sixth order polynomial in $z$. Furthermore, one can directly verify (e.g. using Mathematica) that each of the coefficients in the polynomial is positive for all $p>0,c>0$, giving the result since $z>0$ as well. ∎
@@
giving the result. It therefore suffices perform the following check
  1. (1)
    Fix $10<\lambda<$ 1, $\kappa,p>0$ and let $\ell=\ell(\lambda,\kappa,p_{I})$ as above.
  2. (2)
    If $\ell\leq 1/4$ then $\textrm{CE}^{U}_{0-,\iota}\geq\textrm{CE}^{U}_{0-}$.
  3. (3)
    If $\ell>1/4$ then set $\check{y}=\check{y}(\lambda,\kappa,p_{I})$ and $g=g(\check{y})$ as above. If $g(\check{y})<0$ then $\textrm{CE}^{U}_{0-,\iota}\geq\textrm{CE}^{U}_{0-}$.
This check can easily be performed by any software tool and one always obtains that either $\ell\leq 1/4$ or $g(\check{y})<0$. ∎
@@
We prove this result using the general setup in Appendix A. In this setting, the certainty equivalent absent private information from (33) takes the form
(94) $\textrm{CE}^{i}_{nsn}:=\alpha_{i}\left(\widehat{\Pi}^{\prime}\mu_{X}-\frac{1}{2}\widehat{\Pi}^{\prime}P_{X}^{-1}\widehat{\Pi}\right),\qquad i\in\left\{I,U\right\}.$
Thus, using (2) and (63) we obtain
(95) $\begin{split}\lim_{\alpha_{I}\to 0}\frac{\textrm{CE}^{I}_{nsn}}{\alpha_{I}}&=\frac{1}{\alpha_{U}}\Pi^{\prime}\mu_{X}-\frac{1}{2\alpha_{U}^{2}}\Pi^{\prime}P_{X}^{-1}\Pi;\quad\lim_{\alpha_{I}\to\infty}\textrm{CE}^{I}_{nsn}=\Pi^{\prime}\mu_{X},\\ \lim_{\alpha_{U}\to 0}\textrm{CE}^{I}_{nsn}&=\Pi^{\prime}\mu_{X}-\frac{1}{2\alpha_{I}}\Pi^{\prime}P_{X}^{-1}\Pi;\quad\lim_{\alpha_{U}\to\infty}\alpha_{U}\textrm{CE}^{I}_{nsn}=\alpha_{I}\Pi^{\prime}\mu_{X}.\end{split}$
For the first item, we fix $\alpha_{I}$ and let $\alpha_{U}\to 0$ which corresponds to $\lambda\to 1$ with $\kappa,p_{I}$ fixed. It is easy to see that
$\frac{(1-\lambda)^{2}\kappa p_{I}+(1+p_{I})(\lambda+\kappa p_{I})^{2}}{\kappa(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}}\to\frac{1}{\kappa(1+p_{I})},$
which, in view of (95), gives
$\lim_{\alpha_{U}\to 0}\textrm{CE}^{I}_{0-}=\Pi^{\prime}\mu_{X}-\frac{1}{2\alpha_{I}}\Pi^{\prime}P_{X}^{-1}\Pi+\alpha_{I}\frac{d}{2}\log\left(1+\frac{1}{\kappa(1+p_{I})}\right).$
In particular, the above means that $\textrm{CE}^{I}_{0-}$ remains bounded. As for $\textrm{CE}^{I}_{0-,\iota}$, (18) implies $(1-\lambda)\widehat{y}\to 1+p_{I}$. This gives
$1+\frac{\kappa p_{I}(1+p_{I})+\widehat{y}^{2}}{\kappa(1+p_{I})(1+2\widehat{y})}=\frac{1}{1-\lambda}\left(1-\lambda+\times\frac{(1-\lambda)\kappa p_{I}(1+p_{I})/\widehat{y}+(1-\lambda)\widehat{y}}{\kappa(1+p_{I})\left(1/\widehat{y}+2\right)}\right)\approx\frac{1}{2\kappa(1-\lambda)}.$
Now since $1-\lambda=\alpha_{U}/(\alpha_{I}+\alpha_{U})$, we get that
$\lim_{\alpha_{U}\to 0}\frac{\textrm{CE}^{I}_{0-,\iota}}{-\log(\alpha_{U})}=\alpha_{I}\frac{d}{2},$
which in particular implies that $\lim_{\alpha_{U}\to 0}\textrm{CE}^{I}_{0-,\iota}=\infty$. Now keep $\alpha_{I}$ fixed and let $\alpha_{U}\to\infty$. This corresponds to $\lambda\to 0$ with $\kappa,p_{I}$ fixed. Straight-forward computations show
$\frac{(1-\lambda)^{2}\kappa p_{I}+(1+p_{I})(\lambda+\kappa p_{I})^{2}}{\kappa(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}}\to\frac{p_{I}}{1+\kappa p_{I}(1+p_{I})},$
which, in view of (95),
$\lim_{\alpha_{U}\to\infty}\textrm{CE}^{I}_{0-}=\alpha_{I}\frac{d}{2}\log\left(1+\frac{p_{I}}{1+\kappa p_{I}(1+p_{I})}\right).$
As for $\textrm{CE}^{I}_{0-,\iota}$, (18) implies $\widehat{y}$ converges to the unique positive solution of $y(1+y)=\kappa p_{I}(1+p_{I})$, which is $(1/2)(\sqrt{1+4\kappa p_{I}(1+p_{I})}-1)$ and enforces $(\kappa p_{I}(1+p_{I})+y^{2})/(1+2y)=y$. Therefore,
$\frac{\kappa p_{I}(1+p_{I})+\widehat{y}^{2}}{\kappa(1+p_{I})(1+2\widehat{y})}\to\frac{\sqrt{1+4\kappa p_{I}(1+p_{I})}-1}{2\kappa(1+p_{I})},$
which gives
$\lim_{\alpha_{U}\to\infty}\textrm{CE}^{I}_{0-,\iota}=\alpha_{I}\frac{d}{2}\log\left(1+\frac{\sqrt{1+4\kappa p_{I}(1+p_{I})}-1}{2\kappa(1+p_{I})}\right).$
Hence,
$\begin{split}\lim_{\alpha_{U}\to\infty}\left(\textrm{CE}^{I}_{0-,\iota}-\textrm{CE}^{I}_{0-}\right)&=\frac{1}{2\kappa(1+p_{I})}\left(\sqrt{1+4\kappa p_{I}(1+p_{I})}-1-\frac{2\kappa p_{I}(1+p_{I})}{1+\kappa p_{I}(1+p_{I})}\right),\\ &\,=:\,\frac{1}{2\kappa(1+p_{I})}\times f(\kappa p_{I}(1+p_{I})).\end{split}$
It is easy to see $f(x)>0$ for all $x>0$ and thus the statement follows. This finishes the proof of item (1).
For the second item, let us fix $\alpha_{U}$ and $\alpha_{I}\to 0$. This corresponds to both $\lambda\to 0$ and $\kappa\to 0$. As such, we again will write $\kappa=\alpha_{I}^{2}p_{N}$ and appeal to (18) when analyzing $\widehat{y}$. For $\textrm{CE}^{I}_{0-}$ one can see
@@
we can substitute $\widetilde{y}$ in for $\widehat{y}$ to obtain
$\frac{(1+p_{I})^{2}\left(1+p_{I}-\alpha_{U}^{2}p_{I}p_{N}\right)}{\alpha_{U}^{4}p_{N}}.$
Using the above, limit (96), and the fact that $\lim_{x\to 0}\log(1+C(x)x)/x=C$ if $C(x)\to C$, we get
@@
We now send $\alpha_{I}\to\infty$. This causes both $\lambda\to 1$ and $\kappa\to\infty$ so we will (c.f. (17)) plug in $\kappa=\alpha_{I}^{2}p_{N}$. Doing this, one can verify
$\alpha_{I}^{2}\times\frac{(1-\lambda)^{2}\kappa p_{I}+(1+p_{I})(\lambda+\kappa p_{I})^{2}}{\kappa(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}}\to\frac{1}{p_{N}(1+p_{I})}.$
Since, $\alpha_{I}\log(1+C/\alpha_{I}^{2})\to 0$ as $\alpha_{I}\to\infty$, we get that $\lim_{\alpha_{I}\to\infty}\textrm{CE}^{I}_{0-}=\Pi^{\prime}\mu_{X}$. As for $\textrm{CE}^{I}_{0-,\iota}$, appealing to (18) we see that, by setting $\widehat{y}=\alpha_{I}\widehat{z}$, dividing by $\alpha_{I}^{2}$ and taking the limit that $\widehat{z}$ converges to a solution of
$0=z^{2}\left(1-\frac{\alpha_{U}z}{1+p_{I}}\right)+p_{I}p_{N}\left(\alpha_{U}z+1\right).$
Furthermore, the exact same arguments as in the proof of Proposition 1.5 show there is a unique positive solution. Given that $\widehat{y}/\alpha_{I}\to\widehat{z}$ we find
$\alpha_{I}\times\frac{\kappa p_{I}(1+p_{I})+\widehat{y}^{2}}{\kappa(1+p_{I})(1+2\widehat{y})}\to\frac{p_{I}p_{N}(1+p_{I})+\widehat{z}^{2}}{2p_{N}(1+p_{I})\widehat{z}}.$
This gives that
$\lim_{\alpha_{I}\to\infty}\textrm{CE}^{I}_{0-,\iota}=\Pi^{\prime}\mu_{X}+\frac{d\left(p_{I}p_{N}(1+p_{I})+\widehat{z}^{2}\right)}{4p_{N}(1+p_{I})\widehat{z}}$
since $\alpha_{I}\log(1+C/\alpha_{I})\to C$ as $\alpha_{I}\to\infty$. The above limits finish the proof of the second item.
Lastly, for item (3), we have from (17) that $\alpha_{U}=\alpha_{I}$ yields $\lambda=1/2$. As such, (35) implies $\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}$ is equivalent to
where $\widehat{y}$ solves (18) with $\lambda=1/2$. Note that
$\dot{f}(y)=\frac{2(y(1+y)-\kappa p_{I}(1+p_{I}))}{(1+2y)^{2}},$
so that $f$ is minimized $(0,\infty)$ at $y_{0}=(1/2)\left(\sqrt{1+4\kappa p_{I}(1+p_{I})}-1\right)$, and this value enforces $f(y_{0})=y_{0}$. Therefore, $f(\widehat{y})\geq f(y_{0})=y_{0}$ and hence if
(97) $y_{0}\geq k\quad\Longleftrightarrow\quad\frac{1}{2}\left(\sqrt{1+4\kappa p_{I}(1+p_{I})}-1\right)\geq\frac{(1+p_{I})\left(\kappa p_{I}+(1+p_{I})(1+2\kappa p_{I})^{2}\right)}{(2+p_{I}+2\kappa p_{I}(1+p_{I}))^{2}},$
$\dot{f}(\widehat{y})>0$ and hence $\widehat{y}>y_{0}$. Therefore, $\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}$ is equivalent to $\widehat{y}\geq\check{y}$, which as shown in the proof of Proposition 1.5, is equivalent to $g(\check{y})\geq 0$ for $g$ defined by the cubic function in (18). However, it can easily be checked by any software tool that either (97) holds, or $g(\check{y})\geq 0$. This gives the result.
@@
We finish by stating and proving the version of Proposition 3.7 under the general setup of Appendix A.
Proposition E.1.
Let Assumption A.4 hold. As $p_{I}\to 0$ we obtain the almost sure limits for the insider
@@
$\begin{split}\lim_{p_{I}\to 0}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0}(G,Z_{N})&=\psi_{I,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}+\frac{1}{2}\bigg{\|}P_{X}^{-1/2}\left(\frac{\lambda}{\alpha_{I}}Z_{N}+(1-\lambda)(\psi_{I,0}-\psi_{U,0})\right)\bigg{\|}^{2},\\ \lim_{p_{I}\to 0}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0,\iota}(G,Z_{N})&=\psi_{I,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}\\ &\qquad\qquad+\frac{1}{2(1-\lambda^{2})}\bigg{\|}P_{X}^{-1/2}\left(\frac{\lambda}{\alpha_{I}}Z_{N}+(1-\lambda)(\psi_{I,0}-\psi_{U,0})\right)\bigg{\|}^{2}.\end{split}$
For the uninformed agent we obtain almost surely
@@
$\begin{split}\lim_{p_{I}\to 0}\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0}(H)&=\psi_{U,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}+\frac{1}{2}\bigg{\|}P_{X}^{-1/2}\left(\frac{\lambda}{\alpha_{I}}Z_{N}+\lambda(\psi_{U,0}-\psi_{I,0})\right)\bigg{\|}^{2},\\ \lim_{p_{I}\to 0}\frac{1}{\alpha_{I}}\textrm{CE}^{U}_{0,\iota}(H_{\iota})&=\psi_{U,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}\\ &\qquad\qquad+\frac{1}{2(1-\lambda)^{2}(1+\lambda)^{2}}\bigg{\|}P_{X}^{-1/2}\left(\frac{\lambda}{\alpha_{I}}Z_{N}+(1-\lambda)\lambda(\psi_{U,0}-\psi_{I,0})\right)\bigg{\|}^{2}.\end{split}$
Proof of Proposition E.1.
We will start with the price taking case and use the notation from (17). From Lemma D.1 and using (70) we find
$\begin{split}&\gamma_{I}\textrm{CE}^{I}_{0}(G,Z_{N})=\psi_{I,0}^{\prime}\mu_{X}+\frac{p_{I}(\kappa p_{I}+\lambda)}{(1+p_{I})\left(\frac{1-\lambda}{1+p_{I}}+\kappa p_{I}+\lambda\right)}\psi_{I,0}^{\prime}P_{X}^{-1/2}\bigg{(}\mathcal{E}_{X}+\frac{1}{\sqrt{p_{I}}}\mathcal{E}_{I}+\frac{\mathcal{E}_{N}}{p_{I}\alpha_{I}\sqrt{p_{N}}}\\ &\qquad\qquad-\frac{1+\kappa p_{I}}{p_{I}(\kappa p_{I}+\lambda)}P_{X}^{-1/2}\widehat{\Pi}\bigg{)}+\frac{(1-\lambda)^{2}p_{I}^{2}}{2(1+p_{I})^{3}\left(\frac{1-\lambda}{1+p_{I}}+\kappa p_{I}+\lambda\right)^{2}}\bigg{\|}\mathcal{E}_{X}+\frac{1}{\sqrt{p_{I}}}\mathcal{E}_{I}\\ &\qquad\qquad\qquad-\frac{(1+p_{I})(\kappa p_{I}+\lambda)}{(1-\lambda)p_{I}\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}+\frac{(1+\kappa p_{I})(1+p_{I})}{(1-\lambda)p_{I}}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
It then follows that almost surely
$\begin{split}\lim_{p_{I}\to 0}\gamma_{I}\textrm{CE}^{I}_{0}(G,Z_{N})&=\psi_{I,0}^{\prime}\mu_{X}+\psi_{I,0}^{\prime}P_{X}^{-1/2}\bigg{(}\frac{\lambda}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}-P_{X}^{-1/2}\widehat{\Pi}\bigg{)}\\ &\qquad+\frac{1}{2}\bigg{\|}\frac{\lambda}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}-P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
Substituting back in for $Z_{N}$ in (70), and using (89) gives
$\begin{split}&\psi_{I,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}+\psi_{I,0}^{\prime}P_{X}^{-1}\bigg{(}\frac{\lambda_{I}}{\alpha_{I}}Z_{N}-(1-\lambda)\psi_{U,0}-\lambda\psi_{I,0}\bigg{)}\\ &\qquad\qquad+\frac{1}{2}\bigg{\|}P_{X}^{-1/2}\left(\frac{\lambda}{\alpha_{I}}Z_{N}-(1-\lambda)\psi_{U,0}-\lambda\psi_{I,0}\right)\bigg{\|}^{2}+\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0},\end{split}$
and hence the result. We next consider the price impact case. Here
@@
$\begin{split}&\gamma_{I}\textrm{CE}^{I}_{0,\iota}(G,Z_{N})=\psi_{I,0}^{\prime}\mu_{X}+\psi_{I,0}^{\prime}P_{X}^{-1/2}\bigg{(}\frac{\widehat{y}}{1+p_{I}}\frac{1}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}-\frac{(\widehat{y}+1)^{2}+\kappa p_{I}}{(1-\lambda)(\widehat{y}+1)^{2}}P_{X}^{-1/2}\widehat{\Pi}\\ &\qquad\qquad-\frac{\kappa p_{I}\widehat{y}}{(\widehat{y}+1)^{2}}P_{X}^{-1/2}\psi_{I,0}\bigg{)}+\frac{1}{2(1+p_{I})(1+2\widehat{y})}\bigg{\|}p_{I}\mathcal{E}_{X}+\sqrt{p_{I}}\mathcal{E}_{I}-\frac{\widehat{y}}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}\\ &\qquad\qquad+\frac{(1+p_{I})((\widehat{y}+1)^{2}+\kappa p_{I})}{(1-\lambda)(\widehat{y}+1)^{2}}P_{X}^{-1/2}\widehat{\Pi}+\frac{\widehat{y}\left(\frac{1}{1+p_{I}}(\widehat{y}+1)^{2}+\kappa p_{I}\right)(1+p_{I})}{(\widehat{y}+1)^{2}}P_{X}^{-1/2}\psi_{I,0}\bigg{\|}^{2}.\end{split}$
(18) implies $\widehat{y}\to\lambda/(1-\lambda)$ as $p_{I}\to 0$. It then follows that almost surely
$\begin{split}&\lim_{p_{I}\to 0}\gamma_{I}\textrm{CE}^{I}_{0,\iota}(G,Z_{N})=\psi_{I,0}^{\prime}\mu_{X}+\psi_{I,0}^{\prime}P_{X}^{-1/2}\bigg{(}\frac{\lambda}{(1-\lambda)\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}-\frac{1}{1-\lambda}P_{X}^{-1/2}\widehat{\Pi}\bigg{)}\\ &\qquad\qquad+\frac{1}{2(1-\lambda^{2})}\bigg{\|}\frac{\lambda}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}-P_{X}^{-1/2}\widehat{\Pi}-\lambda P_{X}^{-1/2}\psi_{I,0}\bigg{\|}^{2}.\end{split}$
Substituting back in $Z_{N}$ and plugging in for $\widehat{\Pi}$, the right side is
$\begin{split}&\psi_{I,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}+\psi_{I,0}^{\prime}P_{X}^{-1}\bigg{(}\frac{\lambda}{(1-\lambda)\alpha_{I}}Z_{N}-\psi_{U,0}-\frac{\lambda}{(1-\lambda)}\psi_{I,0}\bigg{)}\\ &\qquad\qquad+\frac{1}{2(1-\lambda^{2})}\bigg{\|}P_{X}^{-1/2}\left(\frac{\lambda}{\alpha_{I}}Z_{N}-(1-\lambda)\psi_{U,0}-2\lambda\psi_{I,0}\right)\bigg{\|}^{2}+\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}.\end{split}$
The result follows by collecting terms. We next turn to $U$ where using (87) we obtain $P_{X}^{1/2}(H-\mu_{X})=\mathcal{E}_{X}+1/\sqrt{p_{I}}\mathcal{E}_{I}+1/(p_{I}\alpha_{I}\sqrt{p_{N}})\mathcal{E}_{N}$. Therefore, from Lemma D.1
@@
$\begin{split}&\gamma_{U}\textrm{CE}^{U}_{0}(H)=\psi_{U,0}^{\prime}\mu_{X}+\frac{(\kappa p_{I}+\lambda)}{(1+p_{I})\left((1-\lambda)\frac{1}{1+p_{I}}+\kappa p_{I}+\lambda\right)}\psi^{\prime}_{U,0}P_{X}^{-1/2}\bigg{(}p_{I}\mathcal{E}_{X}+\sqrt{p_{I}}\mathcal{E}_{I}+\frac{1}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}\\ &\qquad\qquad-\frac{1+\kappa p_{I}}{\kappa p_{I}+\lambda}P_{X}^{-1/2}\widehat{\Pi}\bigg{)}\\ &\qquad+\frac{(1+\kappa p_{I})\lambda^{2}}{2(1+p_{I})^{3}\left(\frac{1}{1+p_{I}}+\kappa p_{I}\right)\left(\frac{1-\lambda}{1+p_{I}}+\kappa p_{I}+\lambda\right)^{2}}\bigg{\|}p_{I}\mathcal{E}_{X}+\sqrt{p_{I}}\mathcal{E}_{I}+\frac{1}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}\\ &\qquad\qquad\qquad-\frac{1+(1+p_{I})\kappa p_{I}}{\lambda}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
This gives the almost sure limit
$\begin{split}&\lim_{p_{I}\to 0}\gamma_{U}\textrm{CE}^{U}_{0}(H)=\psi_{U,0}^{\prime}\mu_{X}+\psi^{\prime}_{U,0}P_{X}^{-1/2}\bigg{(}\frac{\lambda}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}-P_{X}^{-1/2}\widehat{\Pi}\bigg{)}+\frac{1}{2}\bigg{\|}\frac{\lambda}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}-P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
Substituting back in $Z_{N}$ and plugging in for $\widehat{\Pi}$, the right side is
$\begin{split}&\psi_{U,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}+\psi^{\prime}_{U,0}P_{X}^{-1}\bigg{(}\frac{\lambda}{\alpha_{I}}Z_{N}-(1-\lambda)\psi_{U,0}-\lambda\psi_{I,0}\bigg{)}\\ &\qquad\qquad+\frac{1}{2}\bigg{\|}P_{X}^{-1/2}\left(\frac{\lambda}{\alpha_{I}}Z_{N}-(1-\lambda)\psi_{U,0}-\lambda\psi_{I,0}\right)\bigg{\|}^{2}+\frac{1}{2}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0},\end{split}$
from which the result can be deduced. Lastly we consider the price impact case where again using (87) we obtain $P_{X}^{1/2}(H_{\iota}-\mu_{X})=\mathcal{E}_{X}+1/\sqrt{p_{I}}\mathcal{E}_{I}+(\widehat{y}+1)/(p_{I}\alpha_{I}\sqrt{p_{N}})\mathcal{E}_{N}$. Therefore, from Lemma D.2

### v2

@@ B.5. Proofs from Section 3
Next define the function (c.f. (27)) as $Q(p,y)=[\kappa p(1+p)+y^{2}]/[\kappa(1+p)(1+2y)]$ so that
@@
At this point, if one plugs these values into the right side of (94) and takes a common denominator, the numerator is a sixth order polynomial in $z$. Furthermore, one can directly verify each of the coefficients in the polynomial is positive for all $p>0,c>0$, giving the result. ∎
@@
giving the result. It therefore suffices perform the following check: (1) Fix $10<\lambda<$ 1, $\kappa,p>0$ and let $\ell=\ell(\lambda,\kappa,p_{I})$ as above; (2) If $\ell\leq 1/4$ then $\textrm{CE}^{U}_{0-,\iota}\geq\textrm{CE}^{U}_{0-}$; (3) If $\ell>1/4$ then set $\check{y}=\check{y}(\lambda,\kappa,p_{I})$ and $g=g(\check{y})$ as above. If $g(\check{y})<0$ then $\textrm{CE}^{U}_{0-,\iota}\geq\textrm{CE}^{U}_{0-}$. This check can easily be performed by any software tool and one always obtains that either $\ell\leq 1/4$ or $g(\check{y})<0$. ∎
@@
We prove this result in the setup in Appendix A. Here, recall the certainty equivalent absent private information (88). Using (1) and (70) we obtain
(95) $\begin{split}\lim_{\alpha_{I}\to 0}\frac{\textrm{CE}^{I}_{nsn}}{\alpha_{I}}&=\frac{1}{\alpha_{U}}\Pi^{\prime}\mu_{X}-\frac{1}{2\alpha_{U}^{2}}\Pi^{\prime}P_{X}^{-1}\Pi.\end{split}$
We first consider when $\alpha_{I}=\alpha_{U}$ so that (see (7)) $\lambda=1/2$. As such, (28) implies $\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}$ is equivalent to
where $\widehat{y}$ solves (6) with $\lambda=1/2$. But $\dot{f}(y)=2(y(1+y)-\kappa p_{I}(1+p_{I}))/(1+2y)^{2}$, hence $f$ is minimized $(0,\infty)$ at $y_{0}=(1/2)(\sqrt{1+4\kappa p_{I}(1+p_{I})}-1)$, and this value enforces $f(y_{0})=y_{0}$. Therefore, $f(\widehat{y})\geq f(y_{0})=y_{0}$ and hence if
(96) $y_{0}\geq k\,\Leftrightarrow\,\frac{1}{2}\left(\sqrt{1+4\kappa p_{I}(1+p_{I})}-1\right)\geq\frac{(1+p_{I})\left(\kappa p_{I}+(1+p_{I})(1+2\kappa p_{I})^{2}\right)}{(2+p_{I}+2\kappa p_{I}(1+p_{I}))^{2}},$
$\dot{f}(\widehat{y})>0$ and hence $\widehat{y}>y_{0}$. Therefore, $\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}$ is equivalent to $\widehat{y}\geq\check{y}$, which as shown in the proof of Proposition 1.1, is equivalent to $g(\check{y})\geq 0$ for $g$ defined by the cubic function in (6). However, it can easily be checked by any software tool that either (96) holds, or $g(\check{y})\geq 0$. This gives the result for $\alpha_{I}=\alpha_{U}$.
We next fix $\alpha_{U}$ and consider $\alpha_{I}\to 0$. This corresponds to both $\lambda\to 0$ and $\kappa\to 0$. As such, we again will write $\kappa=\alpha_{I}^{2}p_{N}$ and appeal to (6) when analyzing $\widehat{y}$. For $\textrm{CE}^{I}_{0-}$ one can see
@@
we can substitute $\widetilde{y}$ in for $\widehat{y}$ to obtain $(1+p_{I})^{2}\left(1+p_{I}-\alpha_{U}^{2}p_{I}p_{N}\right)/\alpha_{U}^{4}p_{N}$. Using the above, limit (97), and the fact that $\lim_{x\to 0}\log[1+C(x)x]/x=C$ if $C(x)\to C$, we get
@@
We finish by proving Proposition 3.6 in the setup of Appendix A, and recalling $\textrm{CE}^{i}_{nsn}$ from (88).
Proposition B.4.
Let Assumption A.4 hold. As $p_{I}\to 0$ we obtain almost surely
@@
$\begin{split}\lim_{p_{I}\to 0}\textrm{CE}^{I}_{0}(G,Z_{N})&=\textrm{CE}^{I}_{nsn}+\frac{\lambda^{2}}{2\alpha_{I}}Z_{N}^{\prime}P_{X}^{-1}Z_{N},\\ \lim_{p_{I}\to 0}\textrm{CE}^{I}_{0,\iota}(G,Z_{N})&=\textrm{CE}^{I}_{nsn}+\frac{\lambda^{2}}{2\alpha_{I}(1-\lambda^{2})}Z_{N}^{\prime}P_{X}^{-1}Z_{N}.\end{split}$
For the uninformed trader we obtain almost surely
@@
$\begin{split}\lim_{p_{I}\to 0}\textrm{CE}^{U}_{0}(H)&=\textrm{CE}^{U}_{nsn}+\frac{\alpha_{U}\lambda^{2}}{\alpha_{I}^{2}}Z_{N}^{\prime}P_{X}^{-1}Z_{N},\\ \lim_{p_{I}\to 0}\textrm{CE}^{U}_{0,\iota}(H_{\iota})&=\textrm{CE}^{U}_{nsn}+\frac{\alpha_{U}\lambda^{2}}{2\alpha_{I}^{2}(1-\lambda)^{2}(1+\lambda)^{2}}Z_{N}^{\prime}P_{X}^{-1}Z_{N}.\end{split}$
Proof of Proposition B.4.
From (81), (87), (89), (91) we obtain
@@
$\begin{split}\textrm{CE}^{I}_{0}(G,Z_{N})&=\textrm{CE}^{o}_{I,0}(G)\\ &\qquad+\frac{\alpha_{I}R}{2(\lambda_{I}+\beta+R\lambda_{U})^{2}}\left|\lambda_{U}(1-R)P_{X}^{1/2}(G-p_{{\mathbb{Q}}_{0}})-(\lambda_{I}+\beta)P_{X}^{-1/2}\frac{Z_{N}}{\alpha_{I}}\right|^{2},\\ \textrm{CE}^{I}_{0,\iota}(G,Z_{N})&=\textrm{CE}^{o}_{I,0}(G)+\frac{\alpha_{I}R}{2(1+2\widehat{y})}\left|\frac{1-R}{R}P_{X}^{1/2}(G-p_{{\mathbb{Q}}_{0}})-\widehat{y}P_{X}^{-1/2}\frac{Z_{N}}{\alpha_{I}}\right|^{2},\\ \end{split}$
where
@@
$\begin{split}\textrm{CE}^{o}_{I,0}(G)&=-\frac{\alpha_{I}R}{2}\left(\widehat{\Pi}^{\prime}P_{X}^{-1}\widehat{\Pi}-2\widehat{\Pi}^{\prime}\left(\mu_{X}+\frac{1-R}{R}G\right)\right),\\ (1-R)G&=(1-R)X_{0}+\sqrt{R(1-R)}P_{X}^{-1/2}\mathcal{E}_{I}.\end{split}$
From (6) and (77) we see that $p_{I}\to 0$ implies $\widehat{y}\to\lambda_{I}/\lambda_{U}$, $\beta\to 0$ and $R\to 1$. The results for $\textrm{CE}^{I}_{0}(G_{,}Z_{N})$ and $\textrm{CE}^{I}_{0,\iota}(G,Z_{N})$ follow from (88) as $\lambda=\lambda_{I}$ and $1=\lambda_{I}+\lambda_{U}$. For the uninformed trader we similarly obtain (also using (82), (83))